Karen wants to advertise how many cookie chips are in each Big Chip cookie at her bakery.
She randomly selects a sample of 41 cookies and finds that the number of chocolate chips per cookie in the sample has a mean of 14.9 and a standard deviation of 1.3.
What is the 98% confidence interval for the number of chocolate chips per cookie for Big Chip cookies?
Answer accurate to one decimal.
_______ < > _________

Answer by Boreal(15235)

98% CI half t-interval for df=41 is t(0.99,df=40)*s/sqrt(n)=2.423*1.3/sqrt(41)=0.49
(14.4, 15.4) units chocolate chips or 14.4 < mu < 15.4